OpenStudy (anonymous):

Two surveyors with two-way radios leave the same point at 9:00 A.M., one walking due south at 1 mi/hr and the other due west at 3 mi/hr . How long can they communicate with one another if each radio has a maximum range of 10 miles? (Round to the nearest minute.)

OpenStudy (anonymous):

So I started like this \[(3x)^2+(1x)^2=10^2\]

OpenStudy (anonymous):

I get my answer to \[+/-\sqrt{10}\] but doesn't accept my answer I also tried putting 3.1 and 3.2 minutes, am I doing something wrong in the problem?